OpenStudy (gracejaylene(1)):

Compute the number of vinyl tiles, measuring 6 in. each side, needed to tile a kitchen measuring 20 ft. by 16 ft.

OpenStudy (gracejaylene(1)):

@Qwertty123 @phi

OpenStudy (qwertty123):

20*12=? 16*12=?

OpenStudy (gracejaylene(1)):

20*12=240 16*12=192

OpenStudy (qwertty123):

192*240=?

OpenStudy (gracejaylene(1)):

46,080

OpenStudy (qwertty123):

46080/36=?

OpenStudy (gracejaylene(1)):

1280

OpenStudy (qwertty123):

YUP! bingo that is how many titles XD

OpenStudy (gracejaylene(1)):

;D

OpenStudy (qwertty123):

lol JK 1280 tiles